# PhyBlock: A Progressive Benchmark for Physical Understanding and Planning via 3D Block Assembly

### v1

@@ Abstract
While vision-language models (VLMs) have demonstrated promising capabilities in reasoning and planning for embodied agents, their ability to comprehend physical phenomena, particularly within structured 3D environments, remains severely limited. To close this gap, we introduce PhyBlock, a progressive benchmark designed to assess VLMs on physical understanding and planning through robotic 3D block assembly tasks. PhyBlock integrates a novel four-level cognitive hierarchy assembly task alongside targeted Visual Question Answering (VQA) samples, collectively aimed at evaluating progressive spatial reasoning and fundamental physical comprehension, including object properties, spatial relationships, and holistic scene understanding. PhyBlock includes 2600 block tasks (400 assembly tasks, 2200 VQA tasks) and evaluates models across three key dimensions: partial completion, failure diagnosis, and planning robustness. We benchmark 21 state-of-the-art VLMs, highlighting their strengths and limitations in physically grounded, multi-step planning. Our empirical findings indicate that the performance of VLMs exhibits pronounced limitations in high-level planning and reasoning capabilities, leading to a notable decline in performance for the growing complexity of the tasks. Error analysis reveals persistent difficulties in spatial orientation and dependency reasoning. Surprisingly, chain-of-thought prompting offers minimal improvements, suggesting spatial tasks heavily rely on intuitive model comprehension. We position PhyBlock as a unified testbed to advance embodied reasoning, bridging vision-language understanding and real-world physical problem-solving.
@@ 1 Introduction
Existing benchmarks james2020rlbench ; liu2023libero ; shridhar2020alfred ; feng2025reflective ; zhang2024vlabench still suffer from two critical limitations: 1) Perception dominance without planning capability. Current frameworks primarily emphasize perceptual understanding while neglecting long-horizon planning, resulting in models that excel at single-step reasoning but demonstrate inadequate inference capacity in complex scenarios; 2) Unrealistic physical assumptions. The prevailing assumption of objects existing in idealized states while ignoring real-world physical interactions significantly undermines their practical applicability. Consequently, we lack a rigorous yardstick that couples high-level language reasoning with the dynamic constraints of the physical world, leaving open whether modern VLMs truly understand how objects interact in three dimensions.
@@
Building on PhyBlock, we conduct a comprehensive evaluation of 25 state-of-the-art open-source and closed-source vision-language models (VLMs) achiam2023gpt ; openai2023gpt4v ; bai2023qwen ; yang2024qwen2 ; li2024llava ; lin2023video , covering diverse architectures and scales. Empirical results uncover three consistent trends. First, a steep performance cliff: mean planning $F_{1}$ scores drop by more than half from the simplest to the most challenging assembly tier, and no model maintains high recall once tasks demand long-horizon sequencing or hidden-support reasoning. Second, a perception–reasoning gap: models answer low-level questions about colour or shape with high accuracy, yet accuracy collapses on counterfactual, causal, or affordance queries, mirroring the assembly failures that stem from unmodelled dynamics. Third, a pair of universal error modes: (i) mis-estimated Euler angles that lead to systematically incorrect block poses, and (ii) ignored support dependencies that violate basic stability—together accounting for the majority of mistakes across architectures. Notably, enabling chain-of-thought prompting in larger models leaves these two error modes virtually unchanged, indicating that more text tokens alone cannot compensate for missing physical priors. Collectively, these findings expose a fundamental shortfall in current multimodal pre-training: while today’s VLMs perceive objects well, they still lack the physical insight and sequential reasoning needed for reliable embodied planning. Bridging this gap will require architectures and training regimes that fuse rich visual embeddings with explicit physics reasoning and interactive feedback, charting a path toward truly capable embodied agents.
@@ 2.2 Vision-Language Models
Recent advances in VLMs have expanded multimodal reasoning capabilities, yet their understanding of physical phenomena remains limited. While foundational models like GPT-4 achiam2023gpt and GPT-4V openai2023gpt4v demonstrate strong visual-language alignment, and open-source alternatives vicuna2023 ; touvron2023llama enable broader experimentation, these systems show fundamental gaps in physical comprehension. Specialized architectures like region-level VLMs cai2024vip ; chen2023shikra ; peng2023kosmos improve spatial awareness for object localization, while video-based models lin2023video ; tan2024koala enhance temporal reasoning, yet neither adequately addresses the physical understanding required for structured 3D assembly. Current benchmarks cai2024temporalbench ; fu2024video primarily assess static perception or functional affordances rather than the core physical reasoning capabilities needed for tasks involving object properties, structural stability, and spatial dependencies.
@@ 2.3 Physical Understanding
Physical understanding bobrow1984qualitative ; hespos2016five ; mccloskey1983intuitive have widespread applications spanning visual reasoning lerer2016learning ; yiclevrer , embodied AI agrawal2016learning ; byravan2017se3 , etc. The primary works battaglia2016interaction ; lerer2016learning ; yiclevrer focused on simple scenarios where visual primitives (e.g., spheres, cubes) are restricted to a limited set of interactions. The following-up works chow2025physbench ; zhao2025mmvu ; wang2024compositional extend the scope to more realistic scenes with real-world objects and complex backgrounds. Wang et.al wang2024compositional introduces the 4D scene representation to simultaneously model dynamic properties of objects and multi-object interactions. PhysBench chow2025physbench is introduced to evaluate VLMs’ physical world understanding capability across a more diverse and comprehensive tasks. Another stream of works focus on spatial intelligence, which requires the understanding the objects positions in 3D space and their relationships in-between.
@@ 3.2 AOV-Based Assembly Evaluation
This dual-representation enables more rigorous evaluation by: (1) computing intermediate metrics to quantify partial completion, even in failure cases; and (2) diagnosing failure modes via systematic analysis of dependency violations, such as missing prerequisites or conflicting operations. Details of the AOV-based evaluation algorithm are provided in Appendix B.1.
@@ 3.4 Dataset Construction
Data Augmentation. The Level-4 block assembly scenes are essentially generated by augmenting Level-3 scenes through compositional transformations. Due to evaluation-scale constraints, we provide 250 high-quality augmented scenes. However, leveraging our open-source code interface, users can scale the dataset to millions of samples without compromising quality. This augmentation can be achieved through various transformations, including color combinations, shape variations, scene compositions, and environmental lighting adjustments.
@@ 4 Experiment
Table 1: Results (%) overview. Evaluation of Vision-Based Block Construction Tasks.
Model Level-1 Level-2 Level-3 Level-4 Overall.
@@
Model Object Property Object Relationship Scene Understanding Dynamic Reasoning Overrall.
$SH$ $CO$ $SI$ $NU$ $Avg$ $RP$ $AP$ $RD$ $RR$ $Avg$ $OC$ $LC$ $TC$ $VP$ $Avg$ $CF$ $PD$ $OR$ $AD$ $Avg$ $Avg$
@@
Note: The 16 abbreviations denote different task types, grouped into four categories. SH: Shape, CO: Color, SI: Size, NU: Number. RP: Relative position, AP: Absolute position, RD: Relative dependency, RR: Relative rotation. OC: Object counting, LC: Layer counting, TC: Type counting, VP: Viewpoint. CF: Counterfactual, PD: Predictive, OR: Ordering, AD: Affordance.
@@ 4.2 Experiment Findings
Performance plummets with increasing task complexity. As shown in Table 1, we evaluated a range of state-of-the-art vision-language models across four difficulty levels of vision-based block construction tasks. The results clearly demonstrate that current models struggle significantly as task complexity increases. The best-performing model, Claude 3.7 Sonnet, achieved the highest overall recall (47.15%) and F1 score (47.36%), yet its performance still sharply declined from simpler (Level-1 recall 75.62%, F1 76.78 %) to more complex tasks (Level-4 recall 40.93%, F1 41.82%). This trend was consistently observed across all evaluated models, highlighting their limitations in handling complex spatial reasoning and multi-step planning tasks.
Current models excel at object properties but remain challenged by complex physical inference. We report the evaluation results of various models on our proposed Physical VQA benchmark, as shown in Table 2. Among the evaluated models, GPT-o3 achieves the highest overall accuracy (70.0%), demonstrating strong generalization across diverse physical reasoning tasks. It notably excels in Object Property (e.g., 90.0% in CO), Object Relationship (86.0% in RD), and Scene Understanding (80.0% in LC). Claude-3.5 Sonnet and Gemini-2.0-flash also show competitive performance, particularly in perceptual tasks such as Shape and Color, though their capabilities on reasoning-heavy tasks (e.g., Counterfactual and Affordance) remain more limited. These results demonstrate progress in multimodal models’ physical perception and reasoning, while underscoring the persistent challenges of causal and temporal understanding.
@@
Euler errors dominate across models, highlighting universal spatial-reasoning gaps. Figure 4 categorizes assembly errors into four types: Euler, dependency, shape, and overflow errors. Smaller models ($\leq$7B parameters, e.g., InternVL2.5-1B) show high error rates, especially in Euler and dependency tasks, while larger open-source models reduce errors but retain dependency issues. Commercial models (GPT-4o, Claude) outperform but still struggle with Euler errors. Notably, reasoning-tuned models (GPT-o1, Claude-3.7-thinking) reduce dependency/overflow errors but not Euler mistakes, underscoring spatial reasoning as a key challenge for future work. As an ablation study, we evaluate model performance under two evaluation paradigms and observe a significant performance boost when strict pose alignment constraints are relaxed, highlighting the model’s limitations in universal spatial reasoning (Figure 3).
Chain-of-thought offers negligible benefit. Table 1 indicate that Claude 3.7 Sonnet performs nearly identically under normal inference and with Thinking enabled, with a similar error distribution. This suggests that chain-of-thought reasoning provides little to no benefit on this benchmark. We posit that spatial understanding of block shapes relies more on the model’s intuitive processing rather than the generation of an extensive reasoning chain. In failure cases, the model often misinterprets the number or structure of blocks at the outset, causing errors to propagate throughout the reasoning process and ultimately affecting the final output.
Performance degrades steeply from perception-level tasks to strategic multi-step planning, exposing VLMs’ limits in cross-modal reasoning and temporal integration. Experimental results reveal a typical hierarchical difficulty distribution across task levels, as shown in Figure 3, where Level 1 exhibits the lowest difficulty while Level 4 demonstrates the highest complexity. This progression highlights the limitations of current VLMs in tasks that transition from perceptual understanding to strategic planning. The performance degradation suggests that as tasks evolve from basic perception (e.g., object recognition) to advanced planning (e.g., multi-step reasoning), VLMs encounter challenges in effectively integrating multimodal information and executing systematic cognitive processes. Potential factors include insufficient contextual reasoning capacity, limited cross-modal alignment precision, and inadequate temporal dependency modeling in complex decision-making scenarios.
@@ A.3 Data Format and Structural Representation of Block Scenes
    • "euler": A 3D vector [roll, pitch, yaw] defining the block’s orientation using Euler angles in degrees, following the XYZ convention.
@@ B.1 Assembly Step Evaluation via AOV Network
Input: Ground truth assembly sequence of blocks $GT$,
@@
Output: Evaluation metrics: True Positives ($TP$), False Positives ($FP$), False Negatives ($FN$)

### v2

@@ Abstract
While vision-language models (VLMs) have demonstrated promising capabilities in reasoning and planning for embodied agents, their ability to comprehend physical phenomena, particularly within structured 3D environments, remains severely limited. To close this gap, we introduce PhyBlock, a progressive benchmark designed to assess VLMs on physical understanding and planning through robotic 3D block assembly tasks. PhyBlock integrates a novel four-level cognitive hierarchy assembly task alongside targeted Visual Question Answering (VQA) samples, collectively aimed at evaluating progressive spatial reasoning and fundamental physical comprehension, including object properties, spatial relationships, and holistic scene understanding. PhyBlock includes 2600 block tasks (400 assembly tasks, 2200 VQA tasks) and evaluates models across three key dimensions: partial completion, failure diagnosis, and planning robustness. We benchmark 23 state-of-the-art VLMs, highlighting their strengths and limitations in physically grounded, multi-step planning. Our empirical findings indicate that the performance of VLMs exhibits pronounced limitations in high-level planning and reasoning capabilities, leading to a notable decline in performance for the growing complexity of the tasks. Error analysis reveals persistent difficulties in spatial orientation and dependency reasoning. We position PhyBlock as a unified testbed to advance embodied reasoning, bridging vision-language understanding and real-world physical problem-solving.
@@ 1 Introduction
Existing benchmarks james2020rlbench; liu2023libero; shridhar2020alfred; feng2025reflective; zhang2024vlabench still suffer from two critical limitationsxiang2025aligning: 1) Perception dominance without planning capability. Current frameworks primarily emphasize perceptual understanding while neglecting long-horizon planning, resulting in models that excel at single-step reasoning but demonstrate inadequate inference capacity in complex scenarios; 2) Unrealistic physical assumptions. The prevailing assumption of objects existing in idealized states while ignoring real-world physical interactions significantly undermines their practical applicability. Consequently, we lack a rigorous yardstick that couples high-level language reasoning with the dynamic constraints of the physical world, leaving open whether modern VLMs truly understand how objects interact in three dimensions.
@@
Building on PhyBlock, we conduct a comprehensive evaluation of 23 state-of-the-art open-source and closed-source vision-language models (VLMs) achiam2023gpt; openai2023gpt4v; bai2023qwen; yang2024qwen2; li2024llava; lin2023video, covering diverse architectures and scales. Empirical results uncover three consistent trends. First, a steep performance cliff: mean planning $F_{1}$ scores drop by more than half from the simplest to the most challenging assembly tier, and no model maintains high recall once tasks demand long-horizon sequencing or hidden-support reasoning. Second, a perception–reasoning gap: models answer low-level questions about colour or shape with high accuracy, yet accuracy collapses on counterfactual, causal, or affordance queries, mirroring the assembly failures that stem from unmodelled dynamics. Third, a pair of universal error modes: (i) mis-estimated the orientation of blocks that lead to systematically incorrect block poses, and (ii) ignored support dependencies that violate basic stability—together accounting for the majority of mistakes across architectures. Notably, enabling thinking mode prompting in larger models leaves these two error modes virtually unchanged, indicating that more text tokens alone cannot compensate for missing physical priors. Collectively, these findings expose a fundamental shortfall in current multimodal pre-training: while today’s VLMs perceive objects well, they still lack the physical insight and sequential reasoning needed for reliable embodied planning. Bridging this gap will require architectures and training regimes that fuse rich visual embeddings with explicit physics reasoning and interactive feedback, charting a path toward truly capable embodied agents.
@@ 2.2 Vision-Language Models
Recent advances in VLMs have expanded multimodal reasoning capabilities and enabled their successful application across diverse domainshan2025roomtour3d; hao2025conav; liang2023cornav; lin2025navcot; lin2025evolvenav; cao2025ground; lin2025inframind; cao2025video, yet their understanding of physical phenomena remains limited. While foundational models like GPT-4 achiam2023gpt and GPT-4V openai2023gpt4v demonstrate strong visual-language alignment, and open-source alternatives vicuna2023; touvron2023llama enable broader experimentation, these systems show fundamental gaps in physical comprehension. Specialized architectures like region-level VLMs cai2024vip; chen2023shikra; peng2023kosmos improve spatial awareness for object localization, while video-based models lin2023video; tan2024koala; weng2024longvlm; cao2022locvtp enhance temporal reasoning, yet neither adequately addresses the physical understanding required for structured 3D assembly. Current benchmarks cai2024temporalbench; fu2024video primarily assess static perception or functional affordances rather than the core physical reasoning capabilities needed for tasks involving object properties, structural stability, and spatial dependencies.
@@ 2.3 Physical Understanding
Physical understanding bobrow1984qualitative; hespos2016five; mccloskey1983intuitive has widespread applications spanning visual reasoning lerer2016learning; yiclevrer; cao2024physgame, embodied AI agrawal2016learning; byravan2017se3, etc. The primary works battaglia2016interaction; lerer2016learning; yiclevrer; wang2025wisafocused on simple scenarios where visual primitives (e.g., spheres, cubes) are restricted to a limited set of interactions. The follow-up works chow2025physbench; zhao2025mmvu; wang2024compositional extend the scope to more realistic scenes with real-world objects and complex backgrounds. Wang et.al wang2024compositional introduces the 4D scene representation to simultaneously model dynamic properties of objects and multi-object interactions. PhysBench chow2025physbench is introduced to evaluate VLMs’ physical world understanding capability across a more diverse and comprehensive tasks. Another stream of works focus on spatial intelligence, which requires the understanding the objects positions in 3D space and their relationships in-between.
@@ 3.2 AOV-Based Assembly Evaluation
This dual-representation enables more rigorous evaluation by: (1) computing intermediate metrics to quantify partial completion, even in failure cases; and (2) diagnosing failure modes via systematic analysis of dependency violations, such as missing prerequisites or conflicting operations.
This AOV framework enables: (1) fine-grained assessment via intermediate completion metrics, (2) diagnostic analysis of failure modes based on violated dependencies, and (3) evaluation of planning robustness across valid sequence variations. Details of the AOV-based evaluation algorithm are provided in Appendix B.1.
@@ 3.4 Dataset Construction
Data Augmentation. The Level-4 block assembly scenes are generated by augmenting Level-3 scenes through compositional transformations. Due to the deliberate focus of PhyBlock as an evaluation benchmark rather than a large-scale training corpus, we adopt a carefully balanced dataset size—comprising 400 assembly tasks and 2,200 VQA samples—which we found sufficient to reliably assess the physical reasoning and planning capabilities of modern VLMs without introducing redundant or overlapping samples. Each scene is generated within a high-fidelity physics simulator and manually verified to ensure physical validity and uniqueness, making large-scale expansion computationally expensive and conceptually unnecessary for diagnostic evaluation.
Importantly, our minimalist design—with simple geometric shapes and plain colors—reduces visual confounds and isolates core physical reasoning capabilities, enabling a controlled and cognitively interpretable evaluation. However, leveraging our open-source code interface, users can easily scale the dataset to millions of scenes through automated compositional transformations (e.g., color combinations, shape variations, scene compositions, and lighting adjustments) without compromising quality or physical realism. This design philosophy balances rigorous benchmark fidelity with extensibility for future large-scale studies.
@@ 4 Experiment
Table 1: Results (%) overview. Evaluation of 3D Block Assembly Planning (One-time Full Planning).
@@
RecRec PrecPrec $F_{\textsubscript{1}}$ RecRec PrecPrec $F_{\textsubscript{1}}$ RecRec PrecPrec $F_{\textsubscript{1}}$ RecRec PrecPrec $F_{\textsubscript{1}}$ RecRec PrecPrec $F_{\textsubscript{1}}$
@@
[…]
SHSH COCO SISI NUNU AvgAvg RPRP APAP RDRD RRRR AvgAvg OCOC LCLC TCTC VPVP AvgAvg CFCF PDPD OROR ADAD AvgAvg AvgAvg
@@
Random 24.9 25.2 25.0 24.9 25.0 23.9 23.8 23.5 23.6 23.7 25.6 25.1 25.3 24.8 25.2 27.8 26.9 4.2 27.9 21.7 24.1
Note: The 16 abbreviations denote different task types, grouped into four categories. SH: Shape, CO: Color, SI: Size, NU: Number. RP: Relative position, AP: Absolute position, RD: Relative dependency, $R$: Relative rotation. OC: Object counting, LC: Layer counting, TC: Type counting, VP: Viewpoint. CF: Counterfactual, PD: Predictive, OR: Ordering, AD: Affordance.
@@ 4.1 Setup
Random Baseline.
To contextualize model performance, we include a random baseline that simulates an unskilled agent acting without perception or reasoning. For the 3D Block Assembly Planning task, actions are sampled uniformly at random from the valid block and placement space within the Genesis simulator. These random plans are scored using the same AOV-based precision, recall, and F1 metrics, yielding an overall F1 of about 8.8%, which represents the expected lower bound of purely stochastic assembly. For the Physical Understanding VQA branch, answers are sampled uniformly from all options. The Ordering subtask requires ranking rather than single-choice selection, lead to slightly lower scores, but the overall accuracy remains around 24%.
Human Expert Upper Bound.
To complement our quantitative evaluation of vision-language models, we additionally establish a human expert upper bound for reference. We conducted a controlled study involving 20 adult participants with academic or engineering backgrounds related to embodied AI and robotics. Each participant was presented with 400 representative tasks randomly sampled from our benchmark, covering both the assembly planning and physical understanding VQA branches. Participants were instructed to answer the questions or design task plans based on the same multimodal inputs (i.e., textual instructions and visual scenes) as used for model inference.
The collected responses were scored using the same evaluation metrics as the benchmark. On average, human experts achieved a score of 378.83 out of 400, corresponding to an overall accuracy of approximately 94.7%. This result provides an empirical upper bound for interpreting model performance and highlights the considerable gap that remains between current VLM capabilities and human-level reasoning in physically grounded planning and understanding tasks.
@@ 4.2 Experiment Findings
Performance plummets with increasing task complexity.
As shown in Table 3, we evaluated a range of state-of-the-art vision-language models across four difficulty levels of vision-based block construction tasks. The results clearly demonstrate that current models struggle significantly as task complexity increases. The best-performing model, Claude 3.7 Sonnet, achieved the highest overall recall (47.15%) and F1 score (47.36%), yet its performance still sharply declined from simpler (Level-1 recall 75.62%, F1 76.78 %) to more complex tasks (Level-4 recall 40.93%, F1 41.82%). This trend was consistently observed across all evaluated models, highlighting their limitations in handling complex spatial reasoning and multi-step planning tasks.
Current models excel at object properties but remain challenged by complex physical inference. We report the evaluation results of various models on our proposed Physical VQA benchmark, as shown in Table 2. Among the evaluated models, GPT-o3 achieves the highest overall accuracy (70.0%), demonstrating strong generalization across diverse physical reasoning tasks. It notably excels in Object Property (e.g., 90.0% in CO), Object Relationship (86.0% in RD), and Scene Understanding (80.0% in LC). Claude-3.5 Sonnet and Gemini-2.0-flash also show competitive performance, particularly in perceptual tasks such as Shape and Color, though their capabilities on reasoning-heavy tasks (e.g., Counterfactual and Affordance) remain more limited. These results highlight recent advances in multimodal models’ abilities to perceive and reason about physical properties, while also indicating that complex causal and temporal reasoning remains a challenging frontier.
Incorrect Orientation Chosen dominate across models, highlighting universal spatial-reasoning gaps. Figure 4 categorizes assembly errors into four types: orientation, dependency, shape, and overflow errors. Smaller models ($\leq$7B parameters, e.g., InternVL2.5-1B) show high error rates, especially in orientation and dependency tasks, while larger open-source models reduce errors but retain dependency issues. Commercial models (GPT-4o, Claude) outperform but still struggle with orientation chosen errors. Notably, reasoning-tuned models (GPT-o1, Claude-3.7-thinking) reduce dependency/overflow errors but not orientation chosen mistakes, underscoring spatial reasoning as a key challenge for future work. As an ablation study, we evaluate model performance under two evaluation paradigms and observe a significant performance boost when strict pose alignment constraints are relaxed, highlighting the model’s limitations in universal spatial reasoning (Figure 3). Originally, we required models to directly predict absolute orientation or relative spatial coordinates, but this setting yielded near-zero success rates—even for tasks trivial to humans. After simplifying the formulation to high-level spatial reasoning, the tasks remained highly challenging.
@@
Thinking Mode offers negligible benefit. Table 3 indicate that Claude 3.7 Sonnet performs nearly identically under normal inference and with thinking mode enabled, with a similar error distribution. This suggests that Thinking Mode reasoning provides little to no benefit on this benchmark. We posit that spatial understanding of block shapes relies more on the model’s intuitive processing rather than the generation of an extensive reasoning chain. In failure cases, the model often misinterprets the number or structure of blocks at the outset, causing errors to propagate throughout the reasoning process and ultimately affecting the final output.
Performance degrades steeply from perception-level tasks to strategic multi-step planning, exposing VLMs’ limits in cross-modal reasoning and temporal integration. Experimental results reveal a typical hierarchical difficulty distribution across task levels, as shown in Figure 3, where Level 1 exhibits the lowest difficulty while Level 4 demonstrates the highest complexity. This progression highlights the limitations of current VLMs in tasks that transition from perceptual understanding to strategic planning. The performance degradation suggests that as tasks evolve from basic perception (e.g., object recognition) to advanced planning (e.g., multi-step reasoning), VLMs encounter challenges in effectively integrating multimodal information and executing systematic cognitive processes. Potential factors include insufficient contextual reasoning capacity, limited cross-modal alignment precision, and inadequate temporal dependency modeling in complex decision-making scenarios. In our experiments, the tasks are organized into four tiers of increasing complexity (Level‑1 through Level‑4). Each successive level introduces additional blocks, more intricate spatial relationships, and higher-order reasoning requirements, thereby posing a progressively greater challenge for vision-language models.
@@ 5 Conclusion
6 Acknowledgments
This work is supported by National Key Research and Development Program of China (2024YFE0203100), National Natural Science Foundation of China (NSFC) under Grants No.62476293, National Postdoctoral Program for Innovative Talents under Grant Number BX20250379, China Postdoctoral Science Foundation under Grant Number 2025M771521, and General Embodied AI Center of Sun Yat-sen University.
@@ A.3 Data Format and Structural Representation of Block Scenes
    • "orientation": A 3D vector [roll, pitch, yaw] defining the block’s orientation using Euler angles in degrees, following the XYZ convention.
@@ B.1 Assembly Step Evaluation via AOV Network
Input: Ground truth assembly sequence of blocks GTGT,
@@
Output: Evaluation metrics: True Positives (TPTP), False Positives (FPFP), False Negatives (FNFN)